OpenStudy (anonymous):

if you run at a speed of 10 miles per hour and cover a distance of 25 miles how long have you been running.. Use the formula t=d/s

OpenStudy (anonymous):

\[t = \frac ds\]\[t = \frac{25}{10}\]\[t = ?

OpenStudy (anonymous):

25/10=2.5

OpenStudy (anonymous):

Yup. 2.5 hours.

OpenStudy (anonymous):

@nishika17 You're not supposed to give out answers by the way.

OpenStudy (anonymous):

But it says how long have you been running so will it be 2.5 minutes or miles @nishika17 @Calcmathlete

OpenStudy (anonymous):

It says miles per hour meaning that it's 2.5 hours.

OpenStudy (anonymous):

yes